OpenStudy (anonymous):

Arithmetic Sequence : If Sn = -441 for the series 19 + 15 + 11 + ... + tn, find n.

OpenStudy (anonymous):

\[\huge \textbf{Welcome To Openstudy}\]

OpenStudy (anonymous):

What is the last term you are given??

mathslover (mathslover):

i think it is n or \(\huge{t_n}\)

OpenStudy (anonymous):

You are asking to whom??

mathslover (mathslover):

Remember : \[\Huge{S_n=\frac{n}{2}{2n+(n-1)d}}\]

OpenStudy (anonymous):

I think the Sum is given and we have to find the n..

OpenStudy (anonymous):

Formula for sum is: \[\huge \color{green}{ S_n = \frac{n}{2}(2a + (n-1)d)}\]

mathslover (mathslover):

Oh yes sorry

mathslover (mathslover):

i put n there

OpenStudy (anonymous):

Here, You have \(S_n = -441\) \(a = 19\) \(d = 15 -19 = -4\) Just plug in these values in the formula and then solve for n...